Tony and Pepper are married, under 65, and are not blind. Tony died in 2019. They have one child, Morgan (under the age of 17), that Pepper continues to provide a home and 100% support for. Tony had W2 earnings of $200,000 for the year, withholding of $35,000. Pepper didn’t have any earnings in 2019; however, she did receive a $2 million insurance policy payout as a result of Tony’s death.

Prior to his death, Tony sold an old business asset of his. At the time of the sale, the accumulated depreciation was $10,000. He had paid $50,000 for it, and he sold it for $12,000. As a result of a disaster (not a presidentially declared one), Pepper had to live in a hotel for 6 months, while her home was being repaired. The hotel cost her $4,500.

She also had the following:

Property taxes of $4,500

Sales taxes of $6,000

Mortgage interest of $5,000

Charitable Contributions of $30,000

Medical expenses of $2,000

Student Loan interest of $6,000 (for Pepper)

What is the amount of Pepper’s refund check or the amount of the check she has to send to the IRS for her 2019 tax return?
Answer #1

Note : The question involves amount to be disclosed by Pepper in her Tax Returns for 2019.

To arrive at this we need to compute Pepper Income for2019:

Sources of income:

Insurance income : 2,000,000

Share of tony earnings 200000-35000=165000

Capital loss of home (50000-10000-12000)= (28000)

   =========

Gross Income 21,37,000

===========

LESS EXPENSES :

1. Hotel Cost 4500

2. Property taxes 4500

3 sales taxes    6000

4 Mortgage interst 5000

5 charitable    30000

6 medical 2000

7 interst on student loan 6000

Total expenses 58000   

NET TAXABLE INCOME = 21,37,000-58,000=20,79,000

Take appropriate tax rate of20% on 20,79,000= 415800 >> amount of check to be prepared for IRS by Pepper.